OpenStudy (anonymous):

Solve for P. R = 2P + 2Q P = 2/(R - Q) P = 2(R - 2Q) P = (R - 2Q)/2

Parth (parthkohli):

R = 2(P + Q) R/2 = P + Q (r/2) - q = p r - 2q = 2p p = (r - 2q)/2 LAST OPTION